Supply voltage=6 V
R1=20
R2=30
Voltage across R1
Voltage across R2
Current through R1
Current through R2
Total current in circuit
1
Expert's answer
2019-04-08T09:28:45-0400

Voltage across R1 equals 6 V.

Voltage across R2 equals 6 V.

Current through R1 equals 6 V/20 Ohm = 0.3 A.

Current through R2 equals 6 V/30 Ohm = 0.2 A.

Total current in circuit equals 0.2 A + 0.3 A = 0.5 A.